Find the product:
(2+ sqrt 6)^2

Answer by stanbon(75887)

Find the product:
(2+ sqrt 6)^2 = (2+sqrt(6))(2+sqrt(6))
----
= 2(2+sqrt(6)) + sqrt(6)(2+sqrt(6))
-----
----
= 4 + 2sqrt(6) + 2sqrt(6) + 6
-----
= 4sqrt(6) + 10
-----
Cheers,
Stan H.
-------------